All That Blooms provides environmentally friendly lawn services for homeowners. Its operating costs are as follows.

Depreciation $ 1,400 per month
Advertising $ 2,745 per month
Insurance $ 1,000 per month
Weed and feed materials $ 19 per lawn
Direct labor $ 15 per lawn
Fuel $ 3 per lawn


All That Blooms charges $ 86  per treatment for the average single-family lawn

Determine the company’s break-even point in number of lawns serviced per month.

Break-even point lawns

Determine the company’s break-even point in dollars. (Round answer to 0 decimal places, e.g. 1,225.)

Break-even point

$
Answer #1

Total variable costs=Weed and feed materials+Direct labor+Fuel

=(19+15+3)=$37 per lawn

Total fixed cost=Depreciation+Advertising+Insurance

=1400+2745+1000=$5145

Contribution margin=Revenue-Variable cost  

=(86-37)=$49 per lawn

Breakeven=Fixed expenses/Contribution margin

=(5145/49)=105 lawns

=(105*86)=$9030